OpenStudy (anonymous):

The maximum speed of the Tupolev Tu-144 airliner is 694 m/s. What is the speed in kilometers per hour?

OpenStudy (kira_yamato):

694 m/s = 694 m / 1 s = (694/1000)/(1/3600) = 694/1000 x 3600 = 2498.4 km/h (QED)

OpenStudy (anonymous):

thanks

OpenStudy (kira_yamato):

NP